List of the Contributors

Kazuo Akutagawa Shizuoka University, smkacta@ipc.shizuoka.ac.jp
Qing-Ming Cheng Josai University, cheng@math.josai.ac.jp
Franki Dillen Katholieke University Leuven, Franki.Dillen@wis.kuleuven.ac.be
Qing Ding Fudan Unversity, qdingk@online.sh.cn
YuXin Dong Fudan University, yxdong@sh.cnuninet.net
Midori Goto Fukuoka Institute of Technology, m-gotou@fit.ac.jp
Jinichi Itoh Kumamoto University, j-itoh@gpo.kumamoto-u.ac.jp
Shigeo Kawai Saga University, kawais@riemann.pd.saga-u.ac.jp
Yoshihisa Kitagawa Utsunomiya University, kitagawa@cc.utsunomiya-u.ac.jp
Masatoshi Kokubu Tokyo Denki University, kokubu@chiba.dendai.ac.jp
Reiko Miyaoka Sophia University, r-miyaok@hoffman.cc.sophia.ac.jp
Katsuhiro Moriya University of Tsukuba, moriya@math.tsukuba.ac.jp
Katsuhiro Nonaka Josai University, ccs79800@syd.odn.ne.jp
Kouei Sekigawa Niigata University, sekigawa@sc.niigata-u.ac.jp
Katsuhiro Shiohama Saga University, shiohama@ms.saga-u.ac.jp
Young Jin Suh Kyungpook National University, yjsuh@bh.knu.ac.kr